OpenStudy (kmullis6):

PLEASE HELP! Am I right? /: -a^2 - 3b^3 + c^2 + 2b^3 - c^2 (-3)^2 - 3(2)^3 + (-3)^2 + 2(2)^3 - (-3)^2 -9 - 24 + -9 + 16 - -9 = -17 Answer choices: a) -67 b) -17 c) -31 d) 1 I'm confused because I calculated line 2 in the calculator, and it came out as 1.

OpenStudy (kmullis6):

a stands for 3 b stands for 2 c stands for -3

OpenStudy (texaschic101):

-3^2 - 3(2^3) + (-3^2) + 2(2^3) - (-3^2) 9 - 3(8) + 9 + 2(8) - (9) 9 - 24 + 9 + 16 - 9 - 33 + 34 1

OpenStudy (the_fizicx99):

-a^2 - 3b^3 + c^2 + 2b^3 - c^2 Replace those values with the given A, B and C \(\ \color{orange}{A}\rightarrow\color{orange}{3} \) \(\ \color{blue}{B} \rightarrow \color{blue}{2}\) \(\ \color{green}{C} \rightarrow \color{green}{-3}\) \(\ \sf \Large -a^2 - 3b^3 + c^2 + 2b^3 - c^2 \Longrightarrow \color{orange}{3}^2 - 3\color{blue}{(2)}^3 + \color{green}{(-3)}^2 + 2\color{blue}{(2)}^3 - \color{green}{(-3)}^2\) \(\ \sf \Large \color{orange}{3}^2 = \color{orange}{(9)} \) \(\ \sf \Large -3\color{blue}{(2)}^3 = -3\color{blue}{(8)} = \color{blue}{(-24)} \) \(\ \sf \Large + \color{green}{(-3)}^2 = \color{green}{(9)} \) \(\ \sf \Large 2\color{blue}{(2)}^3 = 2\color{blue}{(8)} = \color{blue}{(16)} \) \(\ \sf \Large -\color{green}{(-3)}^2 = -\color{green}{(9)} = \color{green}{(-9)}\) \(\ \sf \Large \color{orange}{(9)} + \color{blue}{(-24)} + \color{green}{(9)} + \color{blue}{(16)} + \color{green}{(-9)} = 1 \)

OpenStudy (the_fizicx99):

U_U it cut off... It was suppose to be -(-3)^2 = -(9) = - 9

OpenStudy (kmullis6):

@tHe_FiZiCx99 Isn't it supposed to be -9 tho? -3^2 = -9

OpenStudy (the_fizicx99):

Anything raised to a square is positive regardless of the signs because -3^2 is the same thing as -3 * -3 = 9 2 negatives = positive 1 negative and 1 positive = negative 2 positive = positive

OpenStudy (kmullis6):

@tHe_FiZiCx99 Ok thanks! I don't know how I forgot that. :P

OpenStudy (the_fizicx99):

You're welcome :>